# A composition of simplified physics-based model with neural operator for trajectory-level seismic response predictions of structural systems

### v1

@@ 1 Introduction
Accurate prediction of structural response trajectories under earthquake ground motion excitation is essential for a wide range of engineering applications, including structural design, maintenance planning, damage localization, and digital twin implementation [1, 2, 3, 4, 5]. High-fidelity nonlinear time history analysis remains the primary computational approach for capturing the full complexity of structural dynamics. However, the high computational cost associated with repeated simulations under varying ground motions poses a significant challenge for real-time decision-making, probabilistic risk assessment, and parametric studies [6, 7, 8, 9].
@@
To address this challenge, a variety of surrogate modeling techniques have been developed to emulate nonlinear structural responses at reduced computational cost. Classical approaches include Bayesian linear regression [10], Kriging [11, 12], polynomial chaos expansions [13], and deep neural networks [14, 15]. These models have shown promise in predicting selected engineering demand parameters but typically operate in finite-dimensional vector spaces and do not capture full response trajectories. More recent developments have explored autoregressive formulations [16, 17], state-space models [18, 19], and deep sequence-to-sequence networks [20, 21] for time-history response prediction. While effective in some cases, these approaches often rely on restrictive assumptions such as narrow-band excitations or simplified system dynamics, limiting their applicability across diverse seismic scenarios. Notably, near-fault ground motions and ground motions on stiff soil sites often exhibit wide-band frequency characteristics, restricting the application of conventional dimensionality reduction-based surrogate modeling approaches.
@@
Neural operators offer a promising alternative by learning mappings between infinite-dimensional function spaces rather than finite-dimensional vectors [22, 23, 24, 25]. Designed to approximate nonlinear solution operators of differential equations, i.e., a generalization of the Green’s function, neural operators enable direct regression from input functions (e.g., ground motions) to output functions (e.g., structural response trajectories). Among existing formulations, the Fourier Neural Operator (FNO) [23] introduces a spectral convolution framework that enables efficient, resolution-invariant learning. Other variants include Deep Operator Networks (DeepONet) [22], which separate functional and spatial encodings, and Graph-based Neural Operators (GNO) [26], which support irregular spatial domains. Despite their success in scientific computing, the application of neural operators to nonlinear structural dynamics remains in its early stages.
@@
Recent studies have begun to explore this potential. Goswami et al. [27] proposed a hybrid DeepONet–FNO framework for modeling structural responses under seismic and wind excitations. Cao et al. [28] employed a wavelet DeepONet to predict the dynamic response of floating offshore structures subjected to irregular wave loading. Tainpakdipat et al. [29] utilized the FNO to accelerate large-scale earthquake rupture simulations, while Perrone et al. [30] integrated neural operators with diffusion models to enhance spectral compatibility in synthetic ground motion generation. Nevertheless, most existing efforts adopt fully data-driven formulations without leveraging physical insights derived from structural models. Moreover, performance under diverse ground motion inputs remains largely unexplored.
@@
To enhance generalization and reduce data dependence, physics-guided machine learning has gained increasing attention. Physics-informed neural networks or operators attempt to embed governing partial differential equation (PDE) constraints directly into the loss function as soft penalties [31, 32, 33, 34]. While effective in cases where PDE residuals are available and differentiable, such methods are difficult to train and apply to structural systems governed by complex hysteretic behavior, where explicit residual and gradient forms may be unavailable or intractable.
@@
This study proposes a different strategy: embedding structural dynamics into the learning architecture through a composite operator framework. Specifically, the proposed approach introduces a simplified physics-based simulator/operator between the seismic input and the neural operator. This intermediate step generates a coarse-grained response trajectory using domain-informed approximations such as linearization, modal reduction, or relaxed solvers with coarse time discretization [35, 36, 37, 38, 39]. The neural operator is then trained to correct the discrepancy between this approximation and the full nonlinear response, allowing it to focus on modeling complex dynamics, such as hysteresis and path dependence, that are not captured by the simplified model. This composite architecture preserves the function-to-function learning paradigm while reducing the complexity of the solution operator and improving data efficiency.
@@
While conceptually related strategies have been explored in quantity-of-interest-based surrogate modeling settings [40, 41], their integration within neural operator learning for full trajectory-level structural response prediction remains limited. The present study addresses this gap by introducing a Composite Physics–Neural Operator (C-PhysFNO) framework that unites coarse-grained physics-based approximations with FNO-based operator learning. In addition, a linear regression-based postprocessing step is proposed to further refine predictions and quantify predictive uncertainty. This refinement operates in closed form and incurs negligible computational cost. The proposed framework is validated on three structural systems subjected to both synthetic and recorded ground motions. Numerical results demonstrate that the C-PhysFNO consistently outperforms standard neural operator baselines in both accuracy and data efficiency, particularly in limited-data regimes.
@@ 2.1 Neural operator learning and Fourier formulation
[…]
@@
[…]
@@
where $\circ$ denotes function composition. Neural operator models typically accept structured input in the form of tensors arranged as “Batch–Spatial–Channel,” where the batch axis corresponds to multiple input samples, the spatial axis represents temporal or spatial discretization points, and the channel axis encodes features at each location. In this setting, the operator input is commonly expressed as $[a(\boldsymbol{x}),\boldsymbol{x}]$, indicating that at each location $\boldsymbol{x}$, the input function has value $a(\boldsymbol{x})$. The choice of kernel operator $\mathcal{K}_{l}$ defines different neural operator variants, including the FNO [23], DeepONet [22], and GNO [26].
The FNO replaces the kernel integrals $\mathcal{K}_{l}$ with convolutions defined in the Fourier domain [23]. This formulation enables the model to capture long-range dependencies efficiently and allows for resolution-invariant learning.
@@
Figure 1: Schematic of the FNO architecture [23]. The input function is lifted into a high-dimensional feature space, transformed through multiple spectral convolution layers, and projected back to the output space.
@@ 3.2.1 Equivalent linear system (ELS)
The ELS approach replaces the nonlinear structural model with a linear system characterized by effective stiffness matrix $\mathbf{K}_{\mathrm{eq}}$ [35]. The intermediate trajectory is obtained by solving:
@@
where $\mathbf{M}$ and $\mathbf{C}$ are the mass and damping matrices, and $\mathbf{K}_{\mathrm{eq}}$ can be determined from various linearization approaches [35, 39, 41]. The most straightforward approach is to use the initial elastic stiffness of the structure. Refined approaches include minimizing the mean square error between the linear and nonlinear responses or maximizing the correlations between response quantities of interest. In this work, we adopt the simple approach of using the elastic stiffness. This is because: (i) we did not observe a noticeable performance gain by implementing a more complex ELS scheme in the C-PhysFNO pipeline; and (ii) seismic design codes typically require performing time history analysis on the linear elastic system before applying it to nonlinear systems. Therefore, from an engineering perspective, no additional modeling effort is required.
@@ 3.2.2 Modal reduction
with $\boldsymbol{r}_{r}\in\mathbb{R}^{r}$ denoting the influence vector. This formulation preserves essential nonlinear behavior within a reduced number of degrees of freedom, enabling efficient simulation [36, 38].
@@ 3.2.3 Relaxed solver
The relaxed solver offers a temporal coarse-graining mechanism by solving the full nonlinear system at a coarser time resolution, typically with time steps $\Delta t^{\prime}=k\Delta t$ where $k\in\mathbb{N}$ (e.g., $k=10$ or $20$). The resulting response is linearly interpolated back to the original time grid for compatibility with the training data. This low-resolution simulation captures dominant (slow) dynamics while filtering out high-frequency nonlinear fluctuations. This approximation is particularly useful in systems where high-frequency content is not critical to performance evaluation [37, 40].
@@ 3.3 Neural operator training for structural response correction
The training is framed as a function-to-function regression task over the temporal domain $t\in[0,\tau]$ and across $n_{d}$ structural degrees of freedom. The input to the neural operator is represented as $[\boldsymbol{z}(t),t]$, combining the intermediate surrogate trajectory with the corresponding time coordinate. The architecture of the neural operator adopts the FNO formulation described in Section 2.1, comprising a lifting layer, multiple spectral convolution layers, and a final projection layer.
@@ 5.1.1 Structural model and stochastic excitation
The first example considers a five-story nonlinear shear building model, shown in Figure 3, with restoring forces governed by the Bouc–Wen hysteresis model [42, 43, 41]. The hysteretic force for the $i$-th story is expressed as:
@@ 5.1.2 Prediction performance and discussion
All models are trained using 400 training samples, validated on 100 samples, and evaluated on 200 test samples. Both standard FNO and C-PhysFNO use 6 Fourier kernel layers, 32 Fourier modes, and 64 feature channels. In the standard FNO, the lifting and projection operators are implemented as fully connected layers that transform the input feature channels via the mappings $2\rightarrow 64$ and $64\rightarrow 128\rightarrow n_{d}$ (with $n_{d}=5$ in this example), respectively, where the input dimension $2=1+1$ corresponds to the ground acceleration $a_{g}(t)$ and the time coordinate $t$. For C-PhysFNO, the input channel dimension starts with $n_{d}+1=6$, including the intermediate representation and time, and the network follows the mapping $6\rightarrow 64$ and $64\rightarrow 128\rightarrow n_{d}$. Training is performed using the Adam optimizer with an initial learning rate of 0.001 and a decay factor of 0.5.
Figure 4 presents an example of the predicted displacement trajectories and corresponding absolute errors, i.e., $\epsilon_{i}=|u_{i}-\hat{u}_{i}|$, for a representative test case using C-PhysFNO with the ELS as $\mathcal{H}_{p}$. The results exhibit close agreement with the ground truth nonlinear responses obtained from finite element simulations with errors consistently below 0.01 m. This demonstrates the capability of the proposed model to accurately reconstruct trajectory-level dynamics under wide-band excitations. Figure 5 compares top-story response predictions across all methods under the same test input: the standard FNO, C-PhysFNO with three different coarse-graining strategies, and the linear regression-based refinement applied to the ELS case. The standard FNO fails to capture key response trends, whereas C-PhysFNO accurately reconstructs trajectory-level dynamics across all three $\mathcal{H}_{p}$ types. Postprocessing via linear regression further refines the predictions and quantifies confidence intervals. Notably, the white noise excitation poses a challenging prediction task due to the lack of structure, highlighting the benefit of injecting an intermediate physics-based model.
@@
Figure 5: Comparison of top-story response predictions for the five-story nonlinear system: (a) input excitation, (b) predictions using standard FNO, (c) predictions using C-PhysFNO with three coarse-graining strategies, and (d) predictions from linear regression-refined C-PhysFNO (ELS case). All models are evaluated under the same ground motion input shown in (a). In (d), the shaded area represents the prediction interval given by the posterior mean $\pm$ one standard deviation.
@@
Figure 6 shows the estimated probability distribution of peak displacement during the response time window for the top story, evaluated across the test set. C-PhysFNO and its refinement more accurately match the ground truth distribution compared to standard FNO. Nonetheless, notable discrepancies remain in capturing extreme values. These observations suggest that future research incorporating Bayesian neural operators [44] with active learning could offer improved tail behavior modeling and rare events.
@@
Figure 6: Probability distribution of top-story peak displacement for the five-story nonlinear system.

### v2

@@ 1 Introduction
Accurate prediction of structural response trajectories under earthquake ground motion excitation is essential for a wide range of engineering applications, including structural design, maintenance planning, damage localization, and digital twin implementation padgett2010risk , roy2013fundamental , lin2023digital , zhang2023seismic , jeon2025neural . High-fidelity nonlinear time history analysis remains the primary computational approach for capturing the full complexity of structural dynamics. However, the high computational cost associated with repeated simulations under varying ground motions poses a significant challenge for real-time decision-making, probabilistic risk assessment, and parametric studies moehle2004framework , kim2021quantile , morato2022optimal , kim2025uncertainty .
@@
To address this challenge, a variety of surrogate modeling techniques have been developed to emulate nonlinear structural responses at reduced computational cost. Classical approaches include Bayesian linear regression ghosh2020seismic , Kriging kim2023estimation , yi2025multi , polynomial chaos expansions zhu2023seismic , and deep neural networks kim2020probabilistic , torky2021deep . These models have shown promise in predicting selected engineering demand parameters but typically operate in finite-dimensional vector spaces and do not capture full response trajectories. To overcome this limitation, recent efforts have explored autoregressive formulations schar2024feature , wang2024comparison , state-space models li2018gibbs , cheng2025state , and deep sequence-to-sequence networks du2018time , wang2020general for time-history response prediction. For example, Zhang et al. (2019) zhang2019deep applied long short-term memory (LSTM) networks to model temporal dependencies in nonlinear building systems, while Zhang et al. (2020) zhang2020physics proposed a physics-guided convolutional neural network that incorporates physical constraints to improve prediction accuracy. More recently, He and Zhang (2025) he2025mc developed a multi-channel gated recurrent unit-based model capable of generalizing nonlinear dynamics across varying structures and excitations. Recent works have also explored data-driven trajectory modeling using deep learning frameworks. For instance, Atila and Spence (2025) atila2025metamodeling proposed a physics-informed LSTM model for nonlinear stochastic systems, while Li and Spence (2022) li2022metamodeling introduced a deep surrogate framework for high-dimensional structural dynamics under stochastic excitation. Although these methods have shown promising results, these approaches often rely on restrictive assumptions such as narrow-band excitations or simplified system dynamics, limiting their applicability across diverse seismic scenarios. Notably, near-fault ground motions and ground motions on stiff soil sites often exhibit wide-band frequency characteristics, restricting the application of conventional dimensionality reduction-based surrogate modeling approaches.
@@
Neural operators offer a promising alternative by learning mappings between infinite-dimensional function spaces rather than finite-dimensional vectors lu2021learning , li2020fourier , kovachki2023neural , kontolati2024learning . Designed to approximate nonlinear solution operators of differential equations, i.e., a generalization of the Green’s function, neural operators enable direct regression from input functions (e.g., ground motions) to output functions (e.g., structural response trajectories). Among existing formulations, the Fourier Neural Operator (FNO) li2020fourier introduces a spectral convolution framework that enables efficient, resolution-invariant learning. Other variants include Deep Operator Networks (DeepONet) lu2021learning , which separate functional and spatial encodings, and Graph-based Neural Operators (GNO) li2020neural , which support irregular spatial domains. Despite their success in scientific computing, the application of neural operators to nonlinear structural dynamics remains in its early stages.
@@
Recent studies have begun to explore this potential. Goswami et al. goswami2025neural proposed a hybrid DeepONet–FNO framework for modeling structural responses under seismic and wind excitations. Cao et al. cao2024deep employed a wavelet DeepONet to predict the dynamic response of floating offshore structures subjected to irregular wave loading. Tainpakdipat et al. tainpakdipat2025fourier utilized the FNO to accelerate large-scale earthquake rupture simulations, while Perrone et al. perrone2025integrating integrated neural operators with diffusion models to enhance spectral compatibility in synthetic ground motion generation. Nevertheless, most existing efforts adopt fully data-driven formulations without leveraging physical insights derived from structural models. Moreover, performance under diverse ground motion inputs remains largely unexplored.
@@
To enhance generalization and reduce data dependence, physics-guided machine learning has gained increasing attention. Physics-informed neural networks or operators attempt to embed governing partial differential equation (PDE) constraints directly into the loss function as soft penalties goswami2023physics , li2024physics , azizzadenesheli2024neural , eshaghi2025variational . While effective in cases where PDE residuals are available and differentiable, such methods are difficult to train and apply to structural systems governed by complex hysteretic behavior, where explicit residual and gradient forms may be unavailable or intractable.
@@
This study proposes a different strategy: embedding structural dynamics into the learning architecture through a composite operator framework. Specifically, the proposed approach introduces a simplified physics-based simulator/operator between the seismic input and the neural operator. This intermediate step generates a coarse-grained response trajectory using domain-informed approximations such as linearization, modal reduction, or relaxed solvers with coarse time discretization elishakoff2017sixty , peherstorfer2018survey , kramer2019nonlinear , patsialis2020reduced , wang2024optimized . The neural operator is then trained to correct the discrepancy between this approximation and the full nonlinear response, allowing it to focus on modeling complex dynamics, such as hysteresis and path dependence, that are not captured by the simplified model. This composite architecture preserves the function-to-function learning paradigm while reducing the complexity of the solution operator and improving data efficiency.
@@
While conceptually related strategies have been explored in quantity-of-interest-based surrogate modeling settings xian2024physics , kim2024dimensionality , their integration within neural operator learning for full trajectory-level structural response prediction remains limited. The present study addresses this gap by introducing a Composite Physics–Neural Operator (C-PhysFNO) framework that unites coarse-grained physics-based approximations with FNO-based operator learning. In addition, a linear regression-based postprocessing step is proposed to further refine predictions and quantify predictive uncertainty. This refinement operates in closed form and incurs negligible computational cost. The proposed framework is validated on three structural systems subjected to both synthetic and recorded ground motions. Numerical results demonstrate that the C-PhysFNO consistently outperforms standard neural operator baselines in both accuracy and data efficiency, particularly in limited-data regimes.
This composite learning approach shares conceptual similarity with prior work that augments physics-based models using data-driven corrections huang2023fast . However, it differs in both scope and architectural design. For example, Huang et al. (2023) huang2023fast proposed a neural vector-enhanced numerical solver for accelerating time integration in low-dimensional differential equation systems. In contrast, our framework targets the trajectory-level prediction of nonlinear seismic responses using operator learning. Moreover, by incorporating structured physics-based approximations—such as equivalent linearization and modal reduction—as intermediate representations, the C-PhysFNO model mitigates spectral bias and enhances predictive fidelity.
@@ 2.1 Neural operator learning and Fourier formulation
$v^{(0)}(\boldsymbol{x})$ $=\mathcal{Q}(a(\boldsymbol{x}))\,$$\quad$ $\text{(Lifting layer)},$ (3)
@@
$u(\boldsymbol{x})$ $=\mathcal{P}(v^{(L)}(\boldsymbol{x}))\,$$\quad$ $\text{(Projection layer)},$ (5)
@@
where $\circ$ denotes function composition. Neural operator models typically accept structured input in the form of tensors arranged as “Batch–Spatial–Channel,” where the batch axis corresponds to multiple input samples, the spatial axis represents temporal or spatial discretization points, and the channel axis encodes features at each location. In this setting, the operator input is commonly expressed as $a(\boldsymbol{x}),\boldsymbol{x}]$, indicating that at each location $\boldsymbol{x}$, the input function has value $a(\boldsymbol{x})$. The choice of kernel operator $\mathcal{K}_{l}$ defines different neural operator variants, including the FNO li2020fourier , DeepONet lu2021learning , and GNO li2020neural .
The FNO replaces the kernel integrals $\mathcal{K}_{l}$ with convolutions defined in the Fourier domain li2020fourier . This formulation enables the model to capture long-range dependencies efficiently and allows for resolution-invariant learning.
@@
Figure 1: Schematic of the FNO architecture li2020fourier . The input function is lifted into a high-dimensional feature space, transformed through multiple spectral convolution layers, and projected back to the output space.
@@ 3.1 Overview of the proposed approach
It is noted that neural operators may exhibit spectral bias that favors the learning of low-frequency components over high-frequency ones rahaman2019spectral . In seismic response prediction under broadband excitations, this bias can hinder accurate learning of complex dynamics. The proposed composite framework alleviates this issue by introducing a physics-based surrogate $\mathcal{H}_{p}$ that captures essential frequency content, thereby simplifying the residual function to be learned by the neural operator.
@@ 3.2.1 Equivalent linear system (ELS)
The ELS approach replaces the nonlinear structural model with a linear system characterized by effective stiffness matrix $\mathbf{K}_{\mathrm{eq}}$ elishakoff2017sixty . The intermediate trajectory is obtained by solving:
@@
where $\mathbf{M}$ and $\mathbf{C}$ are the mass and damping matrices, and $\mathbf{K}_{\mathrm{eq}}$ can be determined from various linearization approaches elishakoff2017sixty , wang2024optimized , kim2024dimensionality . The most straightforward approach is to use the initial elastic stiffness of the structure. Refined approaches include minimizing the mean square error between the linear and nonlinear responses or maximizing the correlations between response quantities of interest. In this work, we adopt the simple approach of using the elastic stiffness. This is because: (i) we did not observe a noticeable performance gain by implementing a more complex ELS scheme in the C-PhysFNO pipeline; and (ii) seismic design codes typically require performing time history analysis on the linear elastic system before applying it to nonlinear systems. Therefore, from an engineering perspective, no additional modeling effort is required.
@@ 3.2.2 Modal reduction
with $\boldsymbol{r}_{r}\in\mathbb{R}^{r}$ denoting the influence vector. This formulation preserves essential nonlinear behavior within a reduced number of degrees of freedom, enabling efficient simulation peherstorfer2018survey , patsialis2020reduced .
@@ 3.2.3 Relaxed solver
The relaxed solver offers a temporal coarse-graining mechanism by solving the full nonlinear system at a coarser time resolution, typically with time steps $\Delta t^{\prime}=k\Delta t$ where $k\in\mathbb{N}$ (e.g., $k=10$ or $20$). The resulting response is linearly interpolated back to the original time grid for compatibility with the training data. This low-resolution simulation captures dominant (slow) dynamics while filtering out high-frequency nonlinear fluctuations. This approximation is particularly useful in systems where high-frequency content is not critical to performance evaluation kramer2019nonlinear , xian2024physics .
@@ 3.3 Neural operator training for structural response correction
The training is framed as a function-to-function regression task over the temporal domain $t\in[0,\tau]$ and across $n_{d}$ structural degrees of freedom. The input to the neural operator is represented as $\boldsymbol{z}(t),t]$, combining the intermediate surrogate trajectory with the corresponding time coordinate. The architecture of the neural operator adopts the FNO formulation described in Section 2.1, comprising a lifting layer, multiple spectral convolution layers, and a final projection layer.
@@ 5.1.1 Structural model and stochastic excitation
The first example considers a five-story nonlinear shear building model, shown in Figure 3, with restoring forces governed by the Bouc–Wen hysteresis model song2006generalized , ismail2009hysteresis , kim2024dimensionality . The hysteretic force for the $i$-th story is expressed as:
@@ 5.1.2 Prediction performance and discussion
All models are trained using 400 training samples, validated on 100 samples, and evaluated on 200 test samples. Both standard FNO and C-PhysFNO use 6 Fourier kernel layers, 32 Fourier modes, and 64 feature channels. In the standard FNO, the lifting and projection operators are implemented as fully connected layers that transform the input feature channels via the mappings $2\rightarrow 64$ and $64\rightarrow 128\rightarrow n_{d}$ (with $n_{d}=5$ in this example), respectively, where the input dimension $2=1+1$ corresponds to the ground acceleration $a_{g}(t)$ and the time coordinate $t$. For C-PhysFNO, the input channel dimension starts with $n_{d}+1=6$, including the intermediate representation and time, and the network follows the mapping $6\rightarrow 64$ and $64\rightarrow 128\rightarrow n_{d}$. Training is performed using the Adam optimizer with an initial learning rate of 0.001 and a decay factor of 0.5. Training of the C-PhysFNO model in this example took approximately 12 minutes using an NVIDIA GeForce RTX 4070 Laptop GPU.
Figure 4 presents an example of the predicted displacement trajectories and corresponding absolute errors, i.e., $\epsilon_{i}=|u_{i}-\hat{u}_{i}|$, for a representative test case using C-PhysFNO with the ELS as $\mathcal{H}_{p}$. The results exhibit close agreement with the ground truth nonlinear responses obtained from finite element simulations with errors consistently below 0.01 m. This demonstrates the capability of the proposed model to accurately reconstruct trajectory-level dynamics under wide-band excitations. To further assess how the intermediate representation mitigates spectral bias rahaman2019spectral , we examine the frequency-domain characteristics of the relevant signals. Figure 5 presents the time histories and power spectral densities (PSDs) for: (i) the input ground acceleration $a_{g}(t)$, (ii) the intermediate response $z(t)$ from the ELS model and the true nonlinear response $u(t)$, and (iii) the residual r(t)=u(t)z(t)r(t)=u(t)-z(t), for a representative top-story response. The PSDs are estimated using Welch’s method. While the residual retains some high-frequency components, its spectral content is substantially reduced and more localized compared to the original nonlinear response. This suggests that the residual mapping learned by the FNO is spectrally simpler and statistically better conditioned. By learning to correct the residual rather than the full mapping, the C-PhysFNO reduces the spectral gap between input and output, aligns the encoded inputs with the dominant solution modes, and improves learning efficiency. These results support the hypothesis that physics-informed intermediate representations mitigate spectral bias and enhance the data efficiency and generalization capability of neural operators. It is worth noting that PSD captures second-order statistics of the signal and cannot fully describe nonlinear effects such as hysteresis or rare extreme responses. Therefore, similarity in PSDs between ELS and nonlinear responses does not imply equivalence in trajectory-level dynamics.
@@
Figure 5: Time-domain and frequency-domain analysis of representative top-floor response: (a) Ground acceleration $a_{g}(t)$, (b) intermediate response $z(t)$ and true nonlinear response $u(t)$, (c) residual r(t)=u(t)z(t)r(t)=u(t)-z(t), and (d) PSDs of all signals.
Figure 6 compares top-story response predictions across all methods under the same test input: the standard FNO, C-PhysFNO with three different coarse-graining strategies, and the linear regression-based refinement applied to the ELS case. The standard FNO fails to capture key response trends, whereas C-PhysFNO accurately reconstructs trajectory-level dynamics across all three $\mathcal{H}_{p}$ types. These observations support the hypothesis that broadband stochastic excitations (such as band-limited white noise) lead to highly variable and spectrally complex input–output mappings, which are difficult to learn directly. By introducing a physics-based intermediate representation, C-PhysFNO effectively reduces this spectral gap, simplifying the learning task into a better-conditioned residual regression problem. Postprocessing via linear regression further refines the predictions and quantifies confidence intervals. Notably, the white noise excitation poses a challenging prediction task due to the lack of structure, highlighting the benefit of injecting an intermediate physics-based model.
Figure 6: Comparison of top-story response predictions for the five-story nonlinear system: (a) predictions using standard FNO, (b) predictions using C-PhysFNO with three coarse-graining strategies, and (c) predictions from linear regression-refined C-PhysFNO (ELS case). All models are evaluated under the same ground motion input. In (c), the shaded area represents the prediction interval given by the posterior mean $\pm$ one standard deviation.
@@
Figure 7 presents a scatter plot and probability distribution of the top-story peak displacement during the response time window, evaluated across the test set. C-PhysFNO and its refinement more accurately match the ground truth distribution compared to standard FNO. Nonetheless, notable discrepancies remain in capturing extreme values. This limitation is partly attributed to the nature of the model, which is trained to predict time-domain response trajectories rather than directly estimating peak values. In this example, the broadband excitation introduces high-frequency components and non-stationary behavior, making peak response estimation particularly challenging. Moreover, small misalignments in predicted trajectories can lead to amplified discrepancies in peak statistics. These observations suggest that future research incorporating Bayesian neural operators bulte2025probabilistic with active learning could offer improved tail behavior modeling and rare events.
@@
Figure 7: Top-story peak displacement prediction for the five-story nonlinear system: (a) scatter plot of predicted versus ground truth peak values across the test set, and (b) estimated probability distribution of peak responses.